Let W⊆R

5

W⊆R5 be the set of solutions of the linear homogeneous system given by

x1−x2+4x3−x4−x5=0

−x1+x2−x3+2x4+x5=0

x3−3x4+x5=0

Accordingly,

(i) show that W⊆R5 is a subspace,

(ii) find a basis for W

(iii) dim(W)=?


1
Expert's answer
2022-01-04T10:27:40-0500

Consider the given homogenous system of linear equations.

"\begin{aligned}

x_{1}-x_{2}+4 x_{3}-x_{4}-x_{5} &=0 \\

-x_{1}+x_{2}-x_{3}+2 x_{4}+x_{5} &=0 \\

x_{3}-3 x_{4}+x_{5} &=0

\end{aligned}"

The above system of equations can be expressed as,

"\left[\begin{array}{ccccc}

1 & -1 & 4 & -1 & -1 \\

-1 & 1 & -1 & 2 & 1 \\

0 & 0 & 1 & -3 & 1

\end{array}\right]\left[\begin{array}{l}

x_{2} \\

x_{2} \\

x_{3} \\

x_{4} \\

x_{5}

\end{array}\right]=\left[\begin{array}{l}

0 \\

0 \\

0 \\

0 \\

0

\end{array}\right]
...(1)"

Let W be the set of solutions of the above linear homogeneous system of equations.

"[i] W \subseteq \mathbb{R}^{5}" is a subspace.

clearly "\overline{0}=\left[\begin{array}{l}0 \\ 0 \\ 0 \\ 0 \\ 0\end{array}\right] \in W."

Now, let "x_{1}, x_{2} \in \mathbb{N}."

"\Rightarrow A X_{2}=0 \text { and } A X_{2}=0"

Consider,

"\begin{gathered}

A\left(X_{1}+X_{2}\right)=A X_{1}+A X_{2}=0+0=0 \\

\therefore A\left(X_{1}+X_{2}\right)=0 \Rightarrow X_{1}+X_{2} \in W .

\end{gathered}"

 Thus. vector addition is satisfied.

Further. Let "\lambda \in \mathbb{R}" and "x \in \mathbb{N}."

Now.

 

"\Rightarrow \quad A x=0 \text {. }"

 

 

"A(\lambda x)=\lambda(A x)=\lambda(0)=0"

 

Thus, "A(\lambda x)=0 \Rightarrow \lambda x \in W."

therefore scalar multiplication is satisfied.

Therefore, "\mid N \subseteq \mathbb{R}^{5}" is a subspace.

ii) From (1) we have,

 

"\left[\begin{array}{ccccc}

1 & -1 & 4 & -1 & -1 \\

-1 & 1 & -1 & 2 & 1 \\

0 & 0 & 1 & -3 & 1

\end{array}\right]\left[\begin{array}{l}

x_{2} \\

x_{2} \\

x_{3} \\

x_{4} \\

x_{5}

\end{array}\right]=\left[\begin{array}{l}

0 \\

0 \\

0 \\

0 \\

0

\end{array}\right]"

 

Apply, "R_{2} \rightarrow R_{2}+R_{1}."

 

"\left[\begin{array}{ccccc}

1 & -1 & 4 & -1 & -1 \\

0 & 0 & 3 & 1 & 0 \\

0 & 0 & 1 & -3 & 1

\end{array}\right]\left[\begin{array}{l}

x_{1} \\

x_{2} \\

x_{3} \\

x_{4} \\

x_{5}

\end{array}\right]=\left[\begin{array}{l}

0 \\

0 \\

0 \\

0 \\

0

\end{array}\right]"

 

 

"\left[\begin{array}{ccccc}

R_{3} & \rightarrow & 3 R_{3}-R_{2} & \\

1 & -1 & 4 & -1 & -1 \\

0 & 0 & 3 & 1 & 0 \\

0 & 0 & 0 & -8 & 3

\end{array}\right]\left[\begin{array}{l}

x_{2} \\

x_{2} \\

x_{3} \\

x_{4} \\

x_{5}

\end{array}\right]=\left[\begin{array}{l}

0 \\

0 \\

0 \\

0 \\

0

\end{array}\right]"

 

"\begin{array}{r}

\Rightarrow x_{1}-x_{2}+4 x_{3}-x_{4}-x_{5}=0 \\

3 x_{3}+x_{4}=0 \\

-8 x_{4}+3 x_{5}=0

\end{array}"

 

No. of equations =3

No. of variables =5.

Therefore, No. of free variables =5-3=2.

Let us treat, "x_{2}" and "x_{4}" as free variables.

Thus. we have,

 

"x_{5}=\frac{8}{3} x_{4} \quad x_{3}=\frac{-1}{3} x_{4}"  

And,

 

"\begin{aligned}

x_{1} &=x_{2}-4 x_{3}+x_{4}+x_{5} \\

&=x_{2}+\frac{4}{3} x_{4}+x_{4}+\frac{8}{3} x_{4} \\

x_{1} &=x_{2}+5 x_{4}

\end{aligned}"

 

"\begin{aligned} X=\left[\begin{array}{l}x_{1} \\ x_{2} \\ x_{3} \\ x_{4} \\ x_{5}\end{array}\right] &=\left[\begin{array}{c}x_{2}+5 x_{4} \\ x_{2} \\ -1 / 3 x_{4} \\ x_{4} \\ 8 / 3 x_{4}\end{array}\right]=\left[\begin{array}{l}x_{2} \\ x_{2} \\ 0 \\ 0 \\ 0\end{array}\right]+\left[\begin{array}{c}5 x_{4} \\ 0 \\ -1 / 3 x_{4} \\ x_{4} \\ 8 / 3 x_{4}\end{array}\right]=x_{2}\left[\begin{array}{l}1 \\ 1 \\ 0 \\ 0 \\ 0\end{array}\right]+x_{4}\left[\begin{array}{c}5 \\ 0 \\ -1 / 3 \\ 1 \\ 8 / 3\end{array}\right] \\ & \therefore\left[\begin{array}{c}x_{1} \\ x_{2} \\ x_{3} \\ x_{4} \\ x_{5}\end{array}\right]=x_{2}\left[\begin{array}{c}1 \\ 1 \\ 0 \\ 0 \\ 0\end{array}\right]+x_{4}\left[\begin{array}{c}15 \\ 0 \\ -1 \\ 3 \\ 8\end{array}\right] \text { where } x_{2}, x_{4} \in \mathbb{R} \end{aligned}"

Thus the solution space W is generated by two linearly independent vectors, namely

"\left[\begin{array}{l}

1 \\

1 \\

0 \\

0 \\

0

\end{array}\right] \text { and }\left[\begin{array}{c}

15 \\

0 \\

-1 \\

3 \\

8

\end{array}\right]"

Therefore,

Basis for W is,

"B=\left\{\left[\begin{array}{c}1 \\ 1 \\ 0 \\ 0 \\ 0\end{array}\right],\left[\begin{array}{c}15 \\ 0 \\ -1 \\ 3 \\ 8\end{array}\right]\right\}"

"\underline{or} \quad B=\{(1,1,0,0,0),(15,0,-1,3,8)\} \subseteq \mathbb{R}^{5}"

(iii) Thus, "\operatorname{dim}(W)=2."